# A mixed Petrov–Galerkin Cosserat rod finite element formulation

### v2

@@ 4.5 Constrained cantilever
This experiment shows the capability of the proposed formulation to realize constraints on the deformation. An initially straight rod with length $L=2\pi$ is considered. The rod is clamped at $\xi=0$ with ${\bm{A}}_{IB}(0)={{\bm{I}}}$ and $_{I}{\bm{r}}_{OC}(0)=\mathbf{0}$. Two load cases are considered, where in the first load case only an external force $_{I}{\bm{b}}_{1}=(0,-P,0)$ is applied on the free tip. The scalar force $P$ is given in terms of the bending stiffness around the ${\bm{e}}_{z}^{B}$ axis, $k_{\mathrm{b}_{z}}=2$, and is given by $P=k_{\mathrm{b}_{z}}(\alpha/L)^{2}$, where $\alpha^{2}\in[0,10]$ can be interpreted as scaled load parameter. In the second load case, an additional external moment $_{B}{\bm{c}}_{1}=(0,0,eP)$ with the lever arm $e=2.5$ is applied at the free tip. Note, that both load cases result in a deformation only in the $_{I}{\bm{e}}_{x}^{I}-$$_{I}{\bm{e}}_{y}^{I}-$plane. Furthermore, we consider three different (constrained) constitutive laws for the simulation, resulting in an unconstrained rod, a shear-stiff rod and an inextensible shear-stiff rod. Their inverse stiffness matrices are listed in Table˜5. As the experiment has a planar setup, the chosen values of the shear stiffness in ${\bm{e}}_{z}^{B}$ direction, the torsional stiffness and the bending stiffness around ${\bm{e}}_{y}^{B}$ are not affecting the results. Considering the inextensible shear-stiff rod, also known as Euler’s elastica, the analytical solution can also be found using elliptic integrals as shown in [32, 26, 11]. The rod was discretized using $n_{\mathrm{el}}=4$ of the $\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$ formulation. The load was applied within 40 load increments and the tolerance for Newton’s solver was $\epsilon=10^{-12}$. Figure˜13 shows the centerline curves for different load parameters $\alpha\in\{0,1,2,4,10\}$ as well as the normalized tip placement. The load case with only the force acting is shown in Figure˜13 and the load case with force and moment acting is shown in load in Figure˜13. The results of the inextensible shear-stiff rod are in agreement with the analytical solution, as visible in the middle and the right plots. The results of the unconstrained and shear-stiff rods are in agreements with the results from [32].
@@
$\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$
213.8 455.1 289.1 198.4 383.1 259.5 185.4 355.1 243.0
@@ 4.6 Lateral-torsional buckling of a ribbon
The final experiment validates the proposed formulation in a scenario where many deformation constraints are applied. Following the buckling analysis of [4, 57] of a ribbon, we use the presented formulation to replicate this analysis. Using the cross-section geometry of width $a=0.05$ and thickness $t=0.002$, together with the material properties of Young’s modulus $E=2.1\cdot 10^{11}$ and Poisson’s ratio $\nu=0.4$, the characteristic curvature $\varepsilon^{\star}=[12(1-\nu^{2})]^{-1/2}\frac{t}{a^{2}}$ is computed. We use two different lengths, given by $L_{A}=0.0157/\varepsilon^{\star}$ and $L_{B}=0.315/\varepsilon^{\star}$. The rod is clamped at one side, and initialized such that the length, thickness and width are in ${\bm{e}}_{x}^{I}$, ${\bm{e}}_{y}^{I}$ and ${\bm{e}}_{z}^{I}$ direction, respectively, as shown in Figure˜14. The constrained deformations of the ribbon are extension, both shear deformations and bending around ${\bm{e}}_{y}^{B}$. For the remaining deformations, i.e., torsion $\varepsilon_{t}$ and bending around ${\bm{e}}_{z}^{B}$, $\varepsilon_{b}$, three different material laws are used. They are defined by the strain energy density functions
@@
with the auxiliary function $\varphi$ as defined by [4][Equations 2.8 and B.1], and the auxiliary strain
@@
We refer to $\gamma$ as the dimensionless force and normalize it by division by $\gamma_{\mathrm{crit}}^{\mathrm{rib}}$, which is shown in Figure˜14 and used in the buckling analysis. As the analysis of [4] shows, the solution of is categorized by the three dimensionless parameters $(\nu,L\varepsilon^{\star},\gamma)$. Therefore, the choice of parameters is to some extent arbitrary.
@@
Different configurations of the rod with $W_{\mathrm{rib}}$ potential for both lengths and different loads can be seen in Figure˜14. Figure˜14 shows the bifurcation diagram as the normalized tip displacement in ${\bm{e}}_{y}^{I}$ direction over the normalized load. The figure shows the results of the three different material laws and both lengths of our simulation together with the results from [4]. Our model indicates to the same bifurcation points as the one found by [4] for all material models. Additionally, also the remainder of the normalized displacement diagram is in very good agreement with the one from literature. For the Sadowsky potential, only the straight configurations, obtained for $\gamma<\gamma_{\mathrm{crit}}^{\mathrm{sdw}}$, are using the approximation of $\varepsilon_{\mathrm{sdw}}$ in the second line of ˜31.
@@
(c) Normalized tip displacement in ${\bm{e}}_{y}^{I}$ direction, $\Delta r_{y}={}_{I}{\bm{r}}_{OC}^{\!\mathrm{T}}(1){}_{I}{\bm{e}}_{y}^{I}$ over normalized force. The different material laws $W_{\mathrm{rib}}$, $W_{\mathrm{sdw}}$ and $W_{\mathrm{quad}}$ are indicated by the colors (Refer to caption), (Refer to caption) and (Refer to caption), respectively. The solid lines correspond to the ribbon with length $L_{B}$, the dashed lines correspond to the ribbon with length $L_{A}$. In the zoomed sections, length $L_{A}$ and $L_{B}$ are indicated by crosses and squares respectively. The solutions from [4] are indicated by (Refer to caption).
@@ 5 Conclusion
In this work, we have proposed a total Lagrangian mixed Petrov–Galerkin finite element formulation for Cosserat rods that addresses fundamental numerical challenges of classical rod formulations. The proposed method employs a Lagrange quaternion interpolation that relies on a minimal set of mathematical concepts while avoiding singularities and ensuring objectivity without additional complexity. In contrast to intrinsically locking-free interpolations based on Lie groups, such as SE(3){SE}(3), our approach avoids the need for trigonometric functions or complex tangent mappings. The formulation is based on a Petrov–Galerkin projection of the internal virtual work and employs the Hellinger–Reissner principle, which introduces independent resultant contact forces and moments and thus leads to a truly locking-free finite element formulation. Furthermore, since the same Petrov–Galerkin projection is used, our formulation can be seamlessly combined with other kinematic interpolations that use the same projection, such as the SE(3){SE}(3) interpolation (as exemplified in this paper), the $\mathbb{R}^{12}$ or the $\mathbb{R}^{3}\times{SO}(3)$ interpolation outlined in [24]. With the Petrov–Galerkin method, the expressions for the virtual work contributions become much simpler than with a Bubnov–Galerkin approach, since the interpolation of the virtual rotation is not dependent on the generalized coordinates. Additionally, no Lagrange multiplier is required to ensure unit length of the nodal quaternions. A drawback of the Petrov–Galerkin method is that the iteration matrix for the Newton–Raphson is not symmetric, which might need more computation time when iterative solvers are used.
@@
In this study, we examined the robustness of the mixed formulation in static analyses. Subsequent steps will involve analyzing the robustness of the mixed formulation in dynamic simulations. To facilitate this analysis, the internal virtual work expression in [34] must be replaced by the mixed formulation proposed in this paper. In combination with an appropriate time-integration scheme, we aim to increase robustness and computational efficiency in dynamic simulations to pave the way for real-time simulation applications, which are particularly relevant for control applications, such as model predictive control.
@@ Appendix C Tangent operator for quaternion
Let ${\bm{P}}=(p_{0},{\bm{p}})\in\mathbb{R}^{4}$ with $p_{0}\in\mathbb{R}$ and ${\bm{p}}\in\mathbb{R}^{3}$ be a non-unit quaternion. Following [20, Equation 6.166] with normalization, the orthonormal transformation matrix is defined by
@@
This result is identical to the result given by [58] and can also be found in [67].
@@ D.1 Spatial convergence
the corresponding twist can be obtained by ${\bm{\theta}}=\operatorname{Log}_{{SE}(3)}({\bm{H}}_{\mathcal{I}\mathcal{B}})\in\mathbb{R}^{6}$, where the definition of the SE(3){SE}(3)-logarithm map can be found in [24, Equation 32]. The error measure is given by
@@ D.2 Solver convergence
To quantify the computational performance of the different formulations, the number of iterations and the interim results of the Newton’s method are analyzed. Let the $k$-th Newton-iterate of all unknowns be given by ${\bm{x}}_{[k]}\in\mathbb{R}^{n}$ and let $m$ be the number of required iteration to fulfill the termination criteria $\sqrt{{\bm{f}}({\bm{x}})^{\!\mathrm{T}}{\bm{f}}({\bm{x}})}<\varepsilon\sqrt{n}$. Based on [1], we approximate the local quadratic convergence rate by

### v3

@@ 4.5 Constrained cantilever
This experiment shows the capability of the proposed formulation to realize constraints on the deformation. An initially straight rod with length $L=2\pi$ is considered. The rod is clamped at $\xi=0$ with ${\bm{A}}_{IB}(0)={{\bm{I}}}$ and $_{I}{\bm{r}}_{OC}(0)=\mathbf{0}$. Two load cases are considered, where in the first load case only an external force $_{I}{\bm{b}}_{1}=(0,-P,0)$ is applied on the free tip. The scalar force $P$ is given in terms of the bending stiffness around the ${\bm{e}}_{z}^{B}$ axis, $k_{\mathrm{b}_{z}}=2$, and is given by $P=k_{\mathrm{b}_{z}}(\alpha/L)^{2}$, where $\alpha^{2}\in[0,10]$ can be interpreted as scaled load parameter. In the second load case, an additional external moment $_{B}{\bm{c}}_{1}=(0,0,eP)$ with the lever arm $e=2.5$ is applied at the free tip. Note, that both load cases result in a deformation only in the $_{I}{\bm{e}}_{x}^{I}-$$_{I}{\bm{e}}_{y}^{I}-$plane. Furthermore, we consider three different (constrained) constitutive laws for the simulation, resulting in an unconstrained rod, a shear-stiff rod and an inextensible shear-stiff rod. Their inverse stiffness matrices are listed in Table˜5. As the experiment has a planar setup, the chosen values of the shear stiffness in ${\bm{e}}_{z}^{B}$ direction, the torsional stiffness and the bending stiffness around ${\bm{e}}_{y}^{B}$ are not affecting the results. Considering the inextensible shear-stiff rod, also known as Euler’s elastica, the analytical solution can also be found using elliptic integrals as shown in [undefae, undefy, undefj]. The rod was discretized using $n_{\mathrm{el}}=4$ of the $\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$ formulation. The load was applied within 40 load increments and the tolerance for Newton’s solver was $\epsilon=10^{-12}$. Figure˜13 shows the centerline curves for different load parameters $\alpha\in\{0,1,2,4,10\}$ as well as the normalized tip placement. The load case with only the force acting is shown in Figure˜13 and the load case with force and moment acting is shown in load in Figure˜13. The results of the inextensible shear-stiff rod are in agreement with the analytical solution, as visible in the middle and the right plots. The results of the unconstrained and shear-stiff rods are in agreements with the results from [undefae].
@@
$\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$ 213.8 455.1 289.1 198.4 383.1 259.5 185.4 355.1 243.0
@@ 4.6 Lateral-torsional buckling of a ribbon
The final experiment validates the proposed formulation in a scenario where many deformation constraints are applied. Following the buckling analysis of [undefc, undefaad] of a ribbon, we use the presented formulation to replicate this analysis. Using the cross-section geometry of width $a=0.05$ and thickness $t=0.002$, together with the material properties of Young’s modulus $E=2.1\cdot 10^{11}$ and Poisson’s ratio $\nu=0.4$, the characteristic curvature $\varepsilon^{\star}=[12(1-\nu^{2})]^{-1/2}\frac{t}{a^{2}}$ is computed. We use two different lengths, given by $L_{A}=0.0157/\varepsilon^{\star}$ and $L_{B}=0.315/\varepsilon^{\star}$. The rod is clamped at one side, and initialized such that the length, thickness and width are in ${\bm{e}}_{x}^{I}$, ${\bm{e}}_{y}^{I}$ and ${\bm{e}}_{z}^{I}$ direction, respectively, as shown in Figure˜14. The constrained deformations of the ribbon are extension, both shear deformations and bending around ${\bm{e}}_{y}^{B}$. For the remaining deformations, i.e., torsion $\varepsilon_{t}$ and bending around ${\bm{e}}_{z}^{B}$, $\varepsilon_{b}$, three different material laws are used. They are defined by the strain energy density functions
@@
with the auxiliary function $\varphi$ as defined by [undefc][Equations 2.8 and B.1], and the auxiliary strain
@@
We refer to $\gamma$ as the dimensionless force and normalize it by division by $\gamma_{\mathrm{crit}}^{\mathrm{rib}}$, which is shown in Figure˜14 and used in the buckling analysis. As the analysis of [undefc] shows, the solution of is categorized by the three dimensionless parameters $(\nu,L\varepsilon^{\star},\gamma)$. Therefore, the choice of parameters is to some extent arbitrary.
@@
Different configurations of the rod with $W_{\mathrm{rib}}$ potential for both lengths and different loads can be seen in Figure˜14. Figure˜14 shows the bifurcation diagram as the normalized tip displacement in ${\bm{e}}_{y}^{I}$ direction over the normalized load. The figure shows the results of the three different material laws and both lengths of our simulation together with the results from [undefc]. Our model indicates to the same bifurcation points as the one found by [undefc] for all material models. Additionally, also the remainder of the normalized displacement diagram is in very good agreement with the one from literature. For the Sadowsky potential, only the straight configurations, obtained for $\gamma<\gamma_{\mathrm{crit}}^{\mathrm{sdw}}$, are using the approximation of $\varepsilon_{\mathrm{sdw}}$ in the second line of ˜31.
@@
(c) Normalized tip displacement in ${\bm{e}}_{y}^{I}$ direction, $\Delta r_{y}={}_{I}{\bm{r}}_{OC}^{\!\mathrm{T}}(1){}_{I}{\bm{e}}_{y}^{I}$ over normalized force. The different material laws $W_{\mathrm{rib}}$, $W_{\mathrm{sdw}}$ and $W_{\mathrm{quad}}$ are indicated by the colors (Refer to caption), (Refer to caption) and (Refer to caption), respectively. The solid lines correspond to the ribbon with length $L_{A}$, the dashed lines correspond to the ribbon with length $L_{B}$. In the zoomed sections, length $L_{A}$ and $L_{B}$ are indicated by crosses and squares respectively. The solutions from [undefc] are indicated by (Refer to caption).
@@ 5 Conclusion
In this work, we have proposed a total Lagrangian mixed Petrov–Galerkin finite element formulation for Cosserat rods that addresses fundamental numerical challenges of classical rod formulations. The proposed method employs a Lagrange quaternion interpolation that relies on a minimal set of mathematical concepts while avoiding singularities and ensuring objectivity without additional complexity. In contrast to intrinsically locking-free interpolations based on Lie groups, such as SE(3){SE}(3), our approach avoids the need for trigonometric functions or complex tangent mappings. The formulation is based on a Petrov–Galerkin projection of the internal virtual work and employs the Hellinger–Reissner principle, which introduces independent resultant contact forces and moments and thus leads to a truly locking-free finite element formulation. Furthermore, since the same Petrov–Galerkin projection is used, our formulation can be seamlessly combined with other kinematic interpolations that use the same projection, such as the SE(3){SE}(3) interpolation (as exemplified in this paper), the $\mathbb{R}^{12}$ or the $\mathbb{R}^{3}\times{SO}(3)$ interpolation outlined in [undefw]. With the Petrov–Galerkin method, the expressions for the virtual work contributions become much simpler than with a Bubnov–Galerkin approach, since the interpolation of the virtual rotation is not dependent on the generalized coordinates. Additionally, no Lagrange multiplier is required to ensure unit length of the nodal quaternions. A drawback of the Petrov–Galerkin method is that the iteration matrix for the Newton–Raphson is not symmetric, which might need more computation time when iterative solvers are used.
@@
In this study, we examined the robustness of the mixed formulation in static analyses. Subsequent steps will involve analyzing the robustness of the mixed formulation in dynamic simulations. To facilitate this analysis, the internal virtual work expression in [undefag] must be replaced by the mixed formulation proposed in this paper. In combination with an appropriate time-integration scheme, we aim to increase robustness and computational efficiency in dynamic simulations to pave the way for real-time simulation applications, which are particularly relevant for control applications, such as model predictive control.
@@ Appendix C Tangent operator for quaternion
Let ${\bm{P}}=(p_{0},{\bm{p}})\in\mathbb{R}^{4}$ with $p_{0}\in\mathbb{R}$ and ${\bm{p}}\in\mathbb{R}^{3}$ be a non-unit quaternion. Following [undefs, Equation 6.166] with normalization, the orthonormal transformation matrix is defined by
@@
This result is identical to the result given by [undefaae] and can also be found in [undefaan].
@@ D.1 Spatial convergence
the corresponding twist can be obtained by ${\bm{\theta}}=\operatorname{Log}_{{SE}(3)}({\bm{H}}_{\mathcal{I}\mathcal{B}})\in\mathbb{R}^{6}$, where the definition of the SE(3){SE}(3)-logarithm map can be found in [undefw, Equation 32]. The error measure is given by
@@ D.2 Solver convergence
To quantify the computational performance of the different formulations, the number of iterations and the interim results of the Newton’s method are analyzed. Let the $k$-th Newton-iterate of all unknowns be given by ${\bm{x}}_{[k]}\in\mathbb{R}^{n}$ and let $m$ be the number of required iteration to fulfill the termination criteria $\sqrt{{\bm{f}}({\bm{x}})^{\!\mathrm{T}}{\bm{f}}({\bm{x}})}<\varepsilon\sqrt{n}$. Based on [undef], we approximate the local quadratic convergence rate by